OpenStudy (anonymous):

20 to the power of what = 365 and how do I calculate this? Thanks so much :)

OpenStudy (amistre64):

logs may help

OpenStudy (amistre64):

\[20^n=365\] \[ln(20^n)=ln(365)\] \[n~ln(20)=ln(365)\] \[n=\frac{ln(365)}{ln(20)}\] \[n=log_{20}(365)\]

OpenStudy (anonymous):

Hi thanks for your reply

OpenStudy (amistre64):

youre welcome 20^1 = 20 20^2 = 400 so we know it has to be somewhere between 1 and 2 . but im not sure theres a nice neat way to get there the way it is without a calculator